In a sample of 240 store customers, 72 use visa card. In another sample of 190, 76 used a mastercard. At š›¼ = 0.10, is there a difference in the proportion of people who use each type of credit card?


1
Expert's answer
2021-12-01T14:26:34-0500

The value of the pooled proportion is computed as


"\bar{p}=\dfrac{X_1+X_2}{n_1+n_2}=\dfrac{72+76}{240+190}=0.344186"

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p_1=p_2"

"H_1:p_1\not=p_2"

This corresponds to a two-tailed test, and a z-test for two population proportions will be used.

Based on the information provided, the significance level isĀ "\alpha = 0.10," and the critical value for a two-tailed test isĀ "z_c = 1.6449."

The rejection region for this two-tailed test isĀ "R = \{z: |z| > 1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}_1-\hat{p}_2}{\sqrt{\bar{p}(1-\bar{p})(1/n_1+1/n_2)}}"

"=\dfrac{72/240-76/190}{\sqrt{0.344186(1-0.344186)(1/240+1/190)}}"

"=-2.1675"

Since it is observed thatĀ "|z| = 2.1675> 1.6449= z_c," it is then concluded thatĀ the null hypothesis is rejected.

Using the P-value approach:

The p-value isĀ "p =2P(Z<-2.1675)= 0.0302," and sinceĀ "p= 0.03002<0.10=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population proportionĀ "p_1"Ā is different thanĀ "p_2," at theĀ "\alpha = 0.10" significance level.